OpenStudy (anonymous):

rewrite with a rational exponent? 3sqrt20

OpenStudy (anonymous):

\[3(20)^{1/2}\]

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

I think so I got (20)1/3 with the 1/3 up in right hand corner, I dont know how to do that so I just wanted to clarify

OpenStudy (anonymous):

\[3\sqrt[?]{20?}\]

OpenStudy (anonymous):

sry I didnt mean to put that

OpenStudy (anonymous):

\[\sqrt[3]{20}\], this is the equation? and it ask to rewrite with a rational exponent

OpenStudy (anonymous):

\[\sqrt[n]{a^{m}}=a^{m/n}\] so for example \[\sqrt{10}=10^{1/2}\]remember that 10 is rised to the power 1 another example \[\sqrt[3]{20}=20^{1/3}\]

OpenStudy (anonymous):

so the answer is (20)1/3???

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I see you dont have the parenthesis around it. does it have them or not?

OpenStudy (anonymous):

ever since (20)1/3 = (20)^1/3

OpenStudy (anonymous):

thank u, thank u!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

no problem, I appreciate your help and clarification :)